# EditGRPO: Reinforcement Learning with Post‑Rollout Edits for Clinically Accurate Chest X‑Ray Report Generation

### v1

@@ Abstract
Radiology report generation requires advanced medical image analysis, effective temporal reasoning, and accurate text generation. Although recent innovations, particularly multimodal large language models (MLLMs), have shown improved performance, their supervised fine-tuning (SFT) objective is not explicitly aligned with clinical efficacy. In this work, we introduce EditGRPO, a mixed-policy reinforcement learning (RL) algorithm designed specifically to optimize the generation through clinically motivated rewards. EditGRPO integrates on-policy exploration with off-policy guidance by injecting sentence-level detailed corrections during training rollouts. This mixed-policy approach addresses the exploration dilemma and sampling efficiency issues typically encountered in RL. Applied to a Qwen2.5-VL-3B MLLM initialized with supervised fine-tuning (SFT), EditGRPO outperforms both SFT and vanilla GRPO baselines, achieving an average improvement of 3.4% in CheXbert, GREEN, Radgraph, and RATEScore metrics across four major chest X-ray report generation datasets. Notably, EditGRPO also demonstrates superior out-of-domain generalization, with an average performance gain of 5.9% on unseen datasets.
@@
Kai Zhang1,2thanks: Work done as an intern at NEC Laboratories America., Christopher Malon2, Lichao Sun1, Martin Renqiang Min2 1NEC Laboratories America, 2Lehigh University kaz321@lehigh.edu, malon@nec-labs.com, lis221@lehigh.edu, renqiang@nec-labs.com Code available at: https://github.com/taokz/EditGRPO
@@ 1 Introduction
Automatic generation of chest X-ray reports from medical images represents a significant challenge in multi-modal artificial intelligence (Demner-Fushman et al., 2016; Irvin et al., 2019; Johnson et al., 2019; Zhang et al., 2025b). Effective AI systems in this domain can substantially reduce labor costs and help standardize radiological interpretations. Recent multi-modal large language models (MLLMs), including ChexAgent (Chen et al., 2024) and MAIRA-2 (Bannur et al., 2024), have achieved competitive performance on natural language generation metrics. However, human evaluations indicate these models often hallucinate details or omit critical clinical information (Zhang et al., 2024; Tu et al., 2024; Wu et al., 2024).
To overcome these limitations and improve clinical accuracy, reinforcement learning (RL) presents a promising alternative to supervised fine-tuning (SFT). Unlike SFT, which is sensitive to specific phrasing, RL directly optimizes clinical efficacy metrics, allowing greater flexibility in expression and prioritizing essential clinical content. Previous studies (Zhou et al., 2024b; Yang et al., 2025) have explored Proximal Policy Optimization (PPO) for report generation, but PPO suffers from low sample efficiency due to the need for a value function model. To overcome this drawback, Group Relative Policy Optimization (GRPO) was introduced (Shao et al., 2024), which avoids a value function by computing advantages based on normalized rewards relative to peer samples.
However, GRPO, as an on-policy RL method, inherently depends on the model’s current capabilities, limiting effective skill expansion. Our experiments using QWen2.5-VL-3B (Bai et al., 2025) reveal that this constraint leads to low-quality outputs, frequently defaulting to generic "no finding" statements. Consequently, advantage estimates remain flat, and training performance quickly plateaus, as demonstrated in Fig. 4.
To address this exploration dilemma, we propose EditGRPO, a variant of GRPO that edits generated report candidates during rollout by injecting clinically correct information from reference reports. Edits include deleting or replacing incorrect (false-positive) findings and adding missing (false-negative) findings as shown in Fig. 1. These edits keep the model close to its existing policy while enabling off-policy guidance to bring its generations closer to the reference report.
@@ 2 Method
but Dr. GRPO (Liu et al., 2025b) revealed that vanilla GRPO introduces optimization biases. For example, longer responses are penalized less due to their larger $|o_{i}|$, leading the policy to favor lengthier but potentially incorrect outputs. Additionally, a question-difficulty bias arises from the standard deviation term in the advantage function, where questions with low standard deviation (i.e., very easy or very hard, where rewards are mostly 1 or 0) are disproportionately weighted during policy updates. To mitigate these issues, EditGRPO adopts their proposed modification and removes this bias-inducing quotient as in Dr. GRPO, to use unnormalized advantage. By default, EditGRPO samples $\frac{G}{2}$ responses $o_{i}$, and fills the other half of the batch with edited versions $o^{\prime}_{i}$ of these.
Editing Rule. Given a generated report $\mathbf{x}$ and a reference report $\mathbf{y}$, the RaTE-NER model (Zhao et al., 2024) extracts entities from each, together with vector embeddings and labels indicating presence or absence. Fix a cosine similarity threshold $\tau$. For each reference sentence $y_{j}$, let $E[y_{j}]$ to be the collection of all entities $e$ with $\cos(e,e^{\prime})>\tau$ for some $e^{\prime}\in y_{j}$. An entity $e\in x$ is spurious if it is not in any $E[y_{j}]$ with a matching presence label. RadGraph can also identify edit candidates but is sensitive to phrasing (see Table 4 in the appendix).
@@ 3.1 Experimental Setting
Data. We focus on MIMIC-CXR (Johnson et al., 2019), a well-studied large-scale dataset, for both training and evaluation. Additionally, we incorporate the newly released large-scale RexGradient (Zhang et al., 2025a) dataset to further validate the effectiveness of EditGRPO. We also study two smaller benchmark datasets, IU-XRay (Demner-Fushman et al., 2016) and PadChest-GR (de Castro et al., 2025), both for training and for out-of-domain generalization. See Appendices F and G.
Models. Qwen-VL-2.5 (3B) (Bai et al., 2025) is used for training report generation models. For a comparative baseline on MIMIC-CXR, we include results from other models supporting multi-view and longitudinal inputs, such as ChexAgent (Chen et al., 2024), MAIRA-2 (Bannur et al., 2024), and MedGemma (Sellergren et al., 2025). The performance of the baselines reported in this paper is obtained through our own inference using the curated multi-view and longitudinal datasets. The model’s performance may be influenced by the prompt.
Training. Because of the time and resources required for reinforcement learning, we compare results for MIMIC-CXR at step 1000, corresponding to about 1/4 of the training data, even though performance continues to improve as shown in Fig. 4. For RexGradient, we use the checkpoint at step 600, representing roughly 15% of the data scale. For the smaller datasets, we train for one full epoch. More details are stated in Appendix D and E.
Evaluation. We evaluate the generated reports using radiology-specific metrics, following established protocols. These include RadGraph-F1 (Jain et al., 2021), CheXbert scores (Micro/Macro-F1-14/5) (Smit et al., 2020), RaTE Score (Zhao et al., 2024), and GREEN (Ostmeier et al., 2024). These clinical metrics are designed to emphasize the accuracy of medical findings, with a focus on detecting and assessing clinically relevant entities.
@@ 3.2 Main Results
SFT is necessary before RL for chest X-ray report generation. Table 1 presents detailed results for the MIMIC-CXR dataset, alongside the performance of SFT and EditGRPO methods on RexGradient. GRPO alone achieves an average score of only 17.35%, demonstrating limited effectiveness without prior SFT training. However, incorporating an initial SFT phase significantly boosts performance, yielding approximately a 25% improvement. The same conclusion is also supported by related works (Guo et al., 2025; Fan et al., 2025; Liu et al., 2025a).
EditGRPO surpasses SFT and other RL variants. The SFT+EditGRPO model (or its normalized advantage variant) lead all metrics except RadGraph F1, which is sensitive to the exact wordings prioritized by SFT training. Accounting for semantic equivalence, we see a 5.0% improvement in CheXbert Micro-F1-5, 4.5% improvement in CheXbert Micro-F1-14, and 2.8% improvement in RaTEScore with SFT+EditGRPO. For the macro variants of the CheXbert metrics, the normalized advantage variant of EditGRPO performs slightly better, although its overall performance is generally lower. An ablation where EditGRPO modifies entire $G$ paragraphs rather than individual sentences (limited to one batch element per query) significantly reduces performance. This highlights the importance of small, localized edits closely aligned with the current policy.
@@ 5 Limitation
Although our proposed EditGRPO method demonstrates the robust potential of reinforcement learning to improve the clinical efficacy of chest X-ray report generation, several limitations remain. First, we did not extend training to larger models (e.g., 7B or 32B parameters), primarily due to the multiplicative increase in computational and time requirements. This constraint is especially relevant in our settings involving multi-view and longitudinal data, where multiple image inputs are required and resource efficiency becomes critical. Second, while the current clinical metrics used in evaluation are generally reliable, they may not fully align with expert radiologist assessments. Additional human evaluation will be necessary to more comprehensively validate the quality and safety of reports generated by models trained with EditGRPO. Lastly, this work focuses specifically on chest X-ray report generation. The applicability of EditGRPO to other medical imaging modalities, such as CT or MRI, remains an open question and is a promising direction for future research.
@@ A.1 Radiology Report Generation
Recent research has focused on transformer-based architectures for medical imaging analysis (Li et al., 2023). Some of these systems are differentiated by their ability to support multi-image inputs, including MAIRA-2 (Bannur et al., 2024) and ChexAgent (Chen et al., 2024). The recent leaderboard RexRank (Zhang et al., 2025b) studies only a single-image setting, and the latest leaders such as MedVersa (Zhou et al., 2024a) focus on integrating multiple modules for detection and segmentation and a large-scale data wrangling effort. In contrast, we focus on training algorithm improvements for a commonly used MLLM architecture, rather than an architecture or data contribution.
@@ A.2 Reinforcement Learning
To further enhance clinical accuracy, some methods incorporate RL to optimize for task-specific rewards, such as capturing “clinically relevant” features (Liu et al., 2019; Kaur and Mittal, 2022; Zhou et al., 2024b) or maintaining logical consistency (Delbrouck et al., 2022; Miura et al., 2021). Rule-based reinforcement learning techniques, exemplified by Group Relative Policy Optimization (GRPO), have shown strong potential for large-scale RL applications, particularly in tasks such as mathematical reasoning and code generation, and have recently been extended to multimodal setting (Wei et al., 2025). In the medical domain, GRPO has been explored in visual question answering tasks (Pan et al., 2025; Fan et al., 2025; Lai et al., 2025). However, to the best of our knowledge, it has not yet been applied to radiology report generation.
@@ A.3 Clinical Evaluation Metrics
Evaluating the quality of generated radiology reports is non-trivial. Early works adopted general-domain natural language processing metrics such as ROUGE (Lin, 2004), BLEU (Papineni et al., 2002), and BERTScore (Zhang* et al., 2020). While these metrics are widely used for text evaluation, they treat differences in wording the same as clinically significant errors, failing to reflect medical accuracy. To address this limitation, clinically informed evaluation metrics, such as CheXbert (Smit et al., 2020), RadGraph (Jain et al., 2021), GREEN (Ostmeier et al., 2024), and RaTEScore (Zhao et al., 2024), have been proposed to better assess clinical correctness and utility. CheXbert is based on multi-label classification results for 5 or 13 diseases (along with one extra “normal” label). RadGraph considers literal entity agreement considering the positive or negative context of each entity. GREEN judges recall and precision errors by LLM prompting. RaTEScore is inspired by RadGraph but less sensitive to phrasing by an F1-like computation which allows semantic matching between entities based on a cosine similarity.
@@ Appendix B Group Relative Policy Optimization
Appendix C Motivation on Post-Rollout Edits
A key obstacle to applying reinforcement learning to chest X‑ray report generation is a shortcut bias: because normal studies vastly outnumber abnormal ones, a policy can maximize its expected reward by emitting a generic “No abnormality” statement. This imbalance drives policy collapse, with the model converging to a single high‑probability “normal” mode—an effect that becomes even more pronounced in low‑resource settings (see Table 1, where training with pure GRPO results in low CheXbert Macro-F1 scores.). Unlike classification tasks, a radiology report is a holistic narrative that may describe an arbitrary subset and count of findings, so straightforward class‑frequency re‑weighting (Lin et al., 2017) may not be practical for all findings simultaneously. To counter this collapse we introduce post‑rollout edits: after sampling a candidate report $o_{i}$ from the policy, we inject expert corrections based on the gold-standard report, recompute the edited trajectory’s log‑probability, and use the resulting adjusted advantage to steer the policy gradient toward clinically faithful outputs. To prevent policy drift, we apply this editing at the sentence level,, which also help the model absorb fine‑grained corrections throughout the long‑context report.
Appendix D Reward Design
@@
Most existing MLLM works that utilize the GRPO algorithm aim to incentivize reasoning capabilities in the VQA setting (Lai et al., 2025; Pan et al., 2025; Liu et al., 2025c; Huang et al., 2025; Fan et al., 2025; Liu et al., 2025a). To achieve this, they incorporate a format score that encourages the model to enclose its reasoning process within <think> and </think> tags. However, for report generation, explicit self‑rationalization is unnecessary: the radiology report itself inherently embodies the trackable reasoning of imaging analysis. It performs semantic mapping from image features to clinical description and lays out the inferred relationships among findings and their differential diagnoses. Therefore, we do not incorporate a format reward and instead conduct “no-thinking” GRPO setup.
@@
We adopt a composite reward strategy, as different clinical efficacy metrics evaluate report quality from complementary perspectives. We adopt the following metrics as the primary reward signals without applying weighting, resulting in a composite reward defined as:
We choose Micro-F1 over Macro-F1 because macro-level rewards tend to be too sparse across most datasets, resulting in weak or uninformative learning signals.
While other metrics, such as GREEN, show stronger alignment with human evaluations (Ostmeier et al., 2024), they rely on LLM-based evaluations, which incur significantly higher inference latency and resource costs. The metrics are computed using their official and standardized implementations: RadGraph-F1111https://pypi.org/project/radgraph/0.1.2/, CheXbert-F1222https://pypi.org/project/f1chexbert/, RaTE Score [^3] https://pypi.org/project/RaTEScore/0.5.0/, and GREEN [^4] https://pypi.org/project/green-score/0.0.8/.
@@
Appendix E Training Configurations

### v2

@@ Abstract
Radiology report generation requires advanced medical image analysis, effective temporal reasoning, and accurate text generation. Although recent innovations, particularly multimodal large language models, have shown improved performance, their supervised fine-tuning (SFT) objective is not explicitly aligned with clinical efficacy. In this work, we introduce EditGRPO, a mixed-policy reinforcement learning algorithm designed specifically to optimize the generation through clinically motivated rewards. EditGRPO integrates on-policy exploration with off-policy guidance by injecting sentence-level detailed corrections during training rollouts. This mixed-policy approach addresses the exploration dilemma and sampling efficiency issues typically encountered in RL. Applied to a Qwen2.5-VL-3B, EditGRPO outperforms both SFT and vanilla GRPO baselines, achieving an average improvement of 3.4% in clinical metrics across four major datasets. Notably, EditGRPO also demonstrates superior out-of-domain generalization, with an average performance gain of 5.9% on unseen datasets.
@@
Kai Zhang1,2thanks: Work done as an intern at NEC Laboratories America., Christopher Malon1, Lichao Sun2, Martin Renqiang Min1 1NEC Laboratories America, 2Lehigh University kaz321@lehigh.edu, malon@nec-labs.com, lis221@lehigh.edu, renqiang@nec-labs.com Code available at: https://github.com/taokz/EditGRPO
@@ 1 Introduction
Automatic generation of chest X-ray reports from medical images represents a significant challenge in multi-modal artificial intelligence (iuxray; chexpert; mimiccxr; rexrank). Effective AI systems in this domain can substantially reduce labor costs and help standardize radiological interpretations. Recent multi-modal large language models (MLLMs), including ChexAgent (chexagent) and MAIRA-2 (maira2), have achieved competitive performance on natural language generation metrics. However, human evaluations indicate these models often hallucinate details or omit critical clinical information (zhang2024generalist; tu2024towards; wu2024hallucinationbenchmarkmedicalvisual).
To overcome these limitations and improve clinical accuracy, reinforcement learning (RL) presents a promising alternative to supervised fine-tuning (SFT). Unlike SFT, which is sensitive to specific phrasing, RL directly optimizes clinical efficacy metrics, allowing greater flexibility in expression and prioritizing essential clinical content. Previous studies (zhou2024large; yang2025aligning) have explored Proximal Policy Optimization (PPO) for report generation, but in LLM settings PPO is costly to optimize: it requires an additional value function comparable in size to the policy, and token-level value estimation is difficult under end-of-sequence rewards. To address this, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) removes the value function by computing advantages based on normalized rewards relative to peer samples. However, GRPO, as an on-policy RL method, inherently depends on the model’s current capabilities, limiting effective skill expansion. Our experiments using QWen2.5-VL-3B (bai2025qwen2) reveal that this constraint leads to low-quality outputs, frequently defaulting to generic "no finding" statements. Consequently, advantage estimates remain flat, and training performance quickly plateaus, as demonstrated in Fig. 4.
To address the exploration dilemma, we propose EditGRPO, a variant of GRPO that edits generated report candidates during rollout by injecting clinically correct information from reference reports. Edits include deleting or replacing incorrect (false-positive) findings and adding missing (false-negative) findings as shown in Fig. 1. These edits keep the model close to its existing policy while enabling off-policy guidance to bring its generations closer to the reference report.
@@
2 Motivation for Post-Rollout Edits
A key obstacle to applying reinforcement learning to chest X‑ray report generation is a shortcut bias: because normal studies vastly outnumber abnormal ones, a policy can maximize its expected reward by emitting a generic “No abnormality” statement. This imbalance drives policy collapse, with the model converging to a single high‑probability “normal” mode—an effect that becomes even more pronounced in low‑resource settings (see Table 1, where training with pure GRPO results in low CheXbert Macro-F1 scores.). Unlike classification tasks, a radiology report is a holistic narrative that may describe an arbitrary subset and count of findings, so straightforward class‑frequency re‑weighting (lin2017focal) may not be practical for all findings simultaneously. To counter this collapse we introduce post‑rollout edits: after sampling a candidate report $o_{i}$ from the policy, we inject expert corrections based on the gold-standard report, recompute the edited trajectory’s log‑probability, and use the resulting adjusted advantage to steer the policy gradient toward clinically faithful outputs. Concretely, within each prompt’s group we interleave unedited rollouts with minimally edited counterparts. This creates intrinsic raw-versus-edited contrasts that enlarge group-relative reward gaps on clinically specific findings, yielding a sharper advantage signal while staying close to the current policy. To prevent policy drift, we apply this editing at the sentence level,, which also help the model absorb fine‑grained corrections throughout the long‑context report.
@@ 3 Method
but Dr. GRPO (drgrpo) revealed that vanilla GRPO introduces optimization biases. For example, longer responses are penalized less due to their larger $|o_{i}|$, leading the policy to favor lengthier but potentially incorrect outputs. Additionally, a question-difficulty bias arises from the standard deviation term in the advantage function, where questions with low standard deviation (i.e., very easy or very hard, where rewards are mostly 1 or 0) are disproportionately weighted during policy updates. To mitigate these issues, EditGRPO adopts their proposed modification and removes this bias-inducing quotient as in Dr. GRPO, to use unnormalized advantage. By default, EditGRPO samples $\frac{G}{2}$ responses $o_{i}$, and fills the other half of the batch with edited versions $o^{\prime}_{i}$ of these.
Editing Rule. Given a generated report $\mathbf{x}$ and a reference report $\mathbf{y}$, the RaTE-NER model (zhao-etal-2024-ratescore) extracts entities from each, together with vector embeddings and labels indicating presence or absence. Fix a cosine similarity threshold $\tau$ (see Table 7 for the hyperparameter sensitivity analysis). For each reference sentence $y_{j}$, let $E[y_{j}]$ to be the collection of all entities $e$ with $\cos(e,e^{\prime})>\tau$ for some $e^{\prime}\in y_{j}$. An entity $e\in x$ is spurious if it is not in any $E[y_{j}]$ with a matching presence label. RadGraph can also identify edit candidates but is sensitive to phrasing (see Table 4 in the appendix).
@@ 4.1 Experimental Setting
Data. We focus on MIMIC-CXR (mimiccxr), a well-studied large-scale dataset, for both training and evaluation. Additionally, we incorporate the newly released large-scale RexGradient (zhang2025rexgradient) dataset to further validate the effectiveness of EditGRPO. We also study two smaller benchmark datasets, IU-XRay (iuxray) and PadChest-GR (de2025padchest), both for training and for out-of-domain generalization. See Appendices E and F for more details.
Models. Qwen-VL-2.5 (3B) (bai2025qwen2) is used for training report generation models. For a comparative baseline on MIMIC-CXR, we include results from other models supporting multi-view and longitudinal inputs, such as ChexAgent (chexagent), MAIRA-2 (maira2), and MedGemma (sellergren2025medgemma). The performance of the baselines reported in this paper is obtained through our own inference using the curated multi-view and longitudinal datasets. The model’s performance may be influenced by the prompt.
Training. Because of the time and resources required for reinforcement learning, we compare results for MIMIC-CXR at step 1000, corresponding to about 1/4 of the training data, even though performance continues to improve as shown in Fig. 4. For RexGradient, we use the checkpoint at step 600, representing roughly 15% of the data scale. For the smaller datasets, we train for one full epoch. We adopt a composite reward strategy, as different clinical efficacy metrics evaluate report quality from complementary perspectives.
As primary reward signals, we adopt the following metrics and combine them as an unweighted sum, using Micro-F1 rather than Macro-F1 due to the sparsity and weak supervision of macro-level rewards across most datasets (see Appendices C and D). The resulting composite reward is:
Evaluation. We evaluate the generated reports using radiology-specific metrics, following established protocols. These include RadGraph-F1 (radgraph), CheXbert scores (Micro/Macro-F1-14/5) (smit-etal-2020-combining), RaTE Score (zhao-etal-2024-ratescore), and GREEN (ostmeier2024green). These clinical metrics are designed to emphasize the accuracy of medical findings, with a focus on detecting and assessing clinically relevant entities. We applied a two-sided Wilcoxon signed-rank test, which confirmed statistically significant gains for EditGRPO on most metrics (Table 8).
@@ 4.2 Main Results
SFT is necessary before RL for chest X-ray report generation. Table 1 presents detailed results for the MIMIC-CXR dataset, alongside the performance of SFT and EditGRPO methods on RexGradient. GRPO alone achieves an average score of only 17.35%, demonstrating limited effectiveness without prior SFT training. However, incorporating an initial SFT phase significantly boosts performance, yielding approximately a 25% improvement. The same conclusion is also supported by related works (guo2025deepseek; fan2025chestx; liu2025x).
EditGRPO surpasses SFT and other RL variants. The SFT+EditGRPO model (or its normalized advantage variant) lead all metrics except RadGraph F1, which is sensitive to the exact wordings prioritized by SFT training. Accounting for semantic equivalence, we see a 5.0% improvement in CheXbert Micro-F1-5, 4.5% improvement in CheXbert Micro-F1-14, and 2.8% improvement in RaTEScore with SFT+EditGRPO. For the macro variants of the CheXbert metrics, the normalized advantage variant of EditGRPO performs slightly better, although its overall performance is generally lower. An ablation where EditGRPO modifies the entire paragraphs rather than the individual sentence (limited to one batch element per query) significantly reduces performance. This highlights the importance of small, localized edits closely aligned with the current policy.
@@ 6 Limitation
Although our proposed EditGRPO method demonstrates the robust potential of reinforcement learning to improve the clinical efficacy of chest X-ray report generation, several limitations remain. First, we did not extend training to larger models (e.g., 7B or 32B parameters), primarily due to the multiplicative increase in computational and time requirements. This constraint is especially relevant in our settings involving multi-view and longitudinal data, where multiple image inputs are required and resource efficiency becomes critical. Second, while the current clinical metrics used in evaluation are generally reliable, they may not fully align with expert radiologist assessments. Additional human evaluation will be necessary to more comprehensively validate the quality and safety of reports generated by models trained with EditGRPO. Thirdly, our rule- and NER-based editing prioritizes precision and auditability but may be brittle to paraphrase, clinical severity nuances, and discourse placement. While stronger LLM editors could improve fluency and insertion, they raise hallucination/negation risks and compute cost; a constrained, entity/negation-aware LLM editor is a promising direction for future work. Lastly, this work focuses specifically on chest X-ray report generation. The applicability of EditGRPO to other medical imaging modalities, such as CT or MRI, remains an open question and is a promising direction for future research.
@@ A.1 Radiology Report Generation
Recent research has focused on transformer-based architectures for medical imaging analysis (llavamed). Some of these systems are differentiated by their ability to support multi-image inputs, including MAIRA-2 (maira2) and ChexAgent (chexagent). The recent leaderboard RexRank (rexrank) studies only a single-image setting, and the latest leaders such as MedVersa (medversa) focus on integrating multiple modules for detection and segmentation and a large-scale data wrangling effort. In contrast, we focus on training algorithm improvements for a commonly used MLLM architecture, rather than an architecture or data contribution.
@@ A.2 Reinforcement Learning
To further enhance clinical accuracy, some methods incorporate RL to optimize for task-specific rewards, such as capturing “clinically relevant” features (liu2019clinically; kaur2022cadxreport; zhou2024large) or maintaining logical consistency (delbrouck2022improving; miura2021improving). Rule-based reinforcement learning techniques, exemplified by Group Relative Policy Optimization (GRPO), have shown strong potential for large-scale RL applications, particularly in tasks such as mathematical reasoning and code generation, and have recently been extended to multimodal setting (grpomultimodal). In the medical domain, GRPO has been explored in visual question answering tasks (pan2025medvlm; fan2025chestx; lai2025med). However, to the best of our knowledge, it has not yet been applied to radiology report generation.
@@ A.3 Clinical Evaluation Metrics
Evaluating the quality of generated radiology reports is non-trivial. Early works adopted general-domain natural language processing metrics such as ROUGE (lin-2004-rouge), BLEU (papineni-etal-2002-bleu), and BERTScore (bertscore). While these metrics are widely used for text evaluation, they treat differences in wording the same as clinically significant errors, failing to reflect medical accuracy. To address this limitation, clinically informed evaluation metrics, such as CheXbert (smit-etal-2020-combining), RadGraph (radgraph), GREEN (ostmeier2024green), and RaTEScore (zhao-etal-2024-ratescore), have been proposed to better assess clinical correctness and utility. CheXbert is based on multi-label classification results for 5 or 13 diseases (along with one extra “normal” label). RadGraph considers literal entity agreement considering the positive or negative context of each entity. GREEN judges recall and precision errors by LLM prompting. RaTEScore is inspired by RadGraph but less sensitive to phrasing by an F1-like computation which allows semantic matching between entities based on a cosine similarity.
@@ Appendix B Group Relative Policy Optimization
Appendix C Reward Design
@@
Most existing MLLM works that utilize the GRPO algorithm aim to incentivize reasoning capabilities in the VQA setting (lai2025med; pan2025medvlm; liu2025visual; huang2025vision; fan2025chestx; liu2025x). To achieve this, they incorporate a format score that encourages the model to enclose its reasoning process within <think> and </think> tags. However, for report generation, explicit self‑rationalization is unnecessary: the radiology report itself inherently embodies the trackable reasoning of imaging analysis. It performs semantic mapping from image features to clinical description and lays out the inferred relationships among findings and their differential diagnoses. Therefore, we do not incorporate a format reward and instead conduct “no-thinking” GRPO setup.
@@
While other metrics, such as GREEN, show stronger alignment with human evaluations (ostmeier2024green), they rely on LLM-based evaluations, which incur significantly higher inference latency and resource costs. The metrics are computed using their official and standardized implementations: RadGraph-F1111https://pypi.org/project/radgraph/0.1.2/, CheXbert-F1222https://pypi.org/project/f1chexbert/, RaTE Score [^3] https://pypi.org/project/RaTEScore/0.5.0/, and GREEN [^4] https://pypi.org/project/green-score/0.0.8/.
@@
Appendix D Training Configurations